OpenStudy (anonymous):

5h – 9 = –16 + 6h A) 4 B) –7 C) 7 D) 10 help please :D

OpenStudy (anonymous):

move the 5h to the other side

OpenStudy (anonymous):

5h – 9 = –16 + 6h -5h -5h -9 = -16 + h now you move the 16

OpenStudy (anonymous):

@Hailey! you know how?

OpenStudy (anonymous):

5h-9=-16+6h combine like terms. 5h=-7+6h then combine the variables

OpenStudy (anonymous):

the 16 gets canceled out right :o?

OpenStudy (anonymous):

add 16 to both sides what do you get -9 = -16 + h

OpenStudy (anonymous):

7 = h ?

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

ahhh :D ty i got answer and learned how to do it =D

OpenStudy (anonymous):

B) –7

OpenStudy (anonymous):

@robert2 do give just answers; show us how?

OpenStudy (anonymous):

if h = -7 5h – 9 = -16 + 6h 5(-7) - 9 = -16 + 6*(-7) -35 - 9 = -16 - 42 -44 = -58 is that correct?!